Prove that for any integer n

n, if n

n is an odd integer, then 6n

2

+5n+1

6n2+5n+1 is an even integer.


1
Expert's answer
2021-10-11T06:08:30-0400

Suppose n is odd. Then, "n=2k-1, k\in \N"

"6n^2+5n+1=6(2k-1)^2+5(2k-1)+1\\
=6(4k^2-4k+1)+10k-5+1\\
=24k^2-24k+6+10k-4\\
=24k^2-14k+2\\
=2(12k^2-7k+1)\\
=2M, M=12k^2-7k+1\\
\text{Hence } 6n^2+5n+1 \text{ is even}"